OpenStudy (anonymous):

A cellular phone company offers offers a contract for which the cost C,in dollars, of T minutes of telephoning is given by C=0.25(T-300)+55.95,where it is assumed that T>=300 Minutes. What times will keep cost between 105.70 and 118.95?

OpenStudy (anonymous):

just plug in your desired cost ranges for C and solve for T, the math is very straight forward

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

that is like a construction worker coming up to me during a remodel and trying to say that building a house is easy

OpenStudy (anonymous):

I am bad with math, especially word problems

OpenStudy (anonymous):

ok , ill walk you through this, first we plug in 105.70 for C, so we have the equation 105.7=.25*(T-300)+55.95 subtract the 55.95 from both sides to get 49.75=.25*(T-300) divide by .25 to get T by itself 199=T-300 finally, add 300 to both sides 499=T the second solution is solved identically

OpenStudy (anonymous):

can you tell me if I am right... between 499 and 552 minutes

OpenStudy (anonymous):

if you still need help with the second answer, plug in 118.95 for C and solve just like i did above, so 118.95=.25*(T-300)+55.95 63=.25(T-300) 252=T-300 552=T so your range of times is between 499<=T<= 552 minutes of talk time on the phone to maintain the price range

OpenStudy (anonymous):

SO i WAS RIGHT?

OpenStudy (anonymous):

yup! ^^